# A Comprehensive Survey on Continual Learning in Generative Models

### v3

@@ 2 Overview of Continual Learning
(§ 6.4), color=darkpastelgreen!100, fill=darkpastelgreen!15, very thick, text=black, text width=2.7cm [$\text{L}^{2}\text{DM}$ [138], color=darkpastelgreen!100, fill=darkpastelgreen!15, very thick, text=black, text width=7.5cm] ] ] ]
Figure 3: Taxonomy of Continual Learning in Generative Models.
2.1 Basic Formulation
Continual learning tackles the challenge of acquiring knowledge from dynamically evolving data distributions, where training samples from distinct tasks arrive sequentially. A continual learning model is required to assimilate new knowledge while preserving previously learned information, typically under constraints that limit or prevent access to past data, while maintaining strong performance across all tasks. Formally, suppose there are $T$ continual learning tasks in total. The training samples for task $t$ can be represented as $\mathcal{D}_{t}=\{(x,y)\}$, where $x$ is the input data and $y$ is the corresponding labels. Notably, the types of training data required vary across different generative models. For example, in MLLM, $(x,y)$ typically consists of a triplet comprising input images, user instructions, and target answers $(\textbf{X}_{v},\textbf{X}_{ins},\textbf{X}_{a})$. In VLA models, it can take the form of state-action pairs $(\mathcal{S},\mathcal{A})$, where the state $\mathcal{S}$ includes an image observation and a text instruction. During the training of task $t$, only the data for the current task is accessible. The training objective of each continual learning task can be formalized as follows:
$\mathop{\arg\min}_{\theta}~{}\mathbb{E}_{(x,y)\sim\mathcal{D}_{t}}\left[\ell(f_{\theta}(x),y))\right],$ (1)
2.2 Evaluation Metrics
The evaluation of a model’s continual learning ability typically involves three aspects: overall performance on all learned tasks, the extent of forgetting on previously learned tasks, and the generalization capability to unseen tasks. For clarity, we denote $a_{t,k}$ as the performance on the $k$-th task after the model has been sequentially trained from task 1 to task $t$ [49].
$\text{Avg}_{t}=\frac{1}{t}\sum_{j=1}^{t}\text{Last}_{j},~{}\text{Last}_{t}=\frac{1}{t}\sum_{i=1}^{t}a_{t,i},$ (2)
$f_{k,t}=\max_{i\in\{1,\cdots,t-1\}}(a_{i,k}-a_{t,k}),\forall k<t.$ (3)
$\text{FM}_{t}=\frac{1}{t-1}\sum_{i=1}^{t-1}f_{i,t}.$ (4)
$\text{BWT}_{t}=\frac{1}{t-1}\sum_{k=1}^{t-1}(a_{t,k}-a_{k,k}).$ (5)
The forgetting measure (FM) quantifies catastrophic forgetting, where lower values indicate better retention of previous knowledge. BWT evaluates knowledge transfer between tasks, with positive values (i.e. BWT > 0) reflecting beneficial forward transfer to prior tasks and negative values (i.e. BWT < 0) indicating interference.
Generalization capability can be evaluated by zero-shot transfer (ZT), which is designed to quantify how well zero-shot transferability is maintained throughout the training process. Formally, for a given task $k$, it first calculates the average performance $A_{k}$ of all models trained on preceding tasks when evaluated on task $k$:
$A_{k}=\frac{1}{k-1}\sum_{i=1}^{k-1}a_{i,k},\forall k>1.$ (6)
ZT is the average of $A_{k}$ after training on the final task:
$\text{ZT}=\frac{1}{T-1}\sum_{k=2}^{T}A_{k}.$ (7)
It is worth noting that the calculation of $a_{t,k}$ depends on the specific task type. For example, Bilingual Evaluation Understudy (BLEU) [139] for text generation task, Intersection-over-Union (IoU) [140] for visual grounding task, Fréchet Inception Distance (FID) [141] for text-to-image generation task, and so on. There are several other evaluation criteria available, and we refer readers to their original papers for detailed specifications.
2.3 Taxonomy
@@
Inspired by the collaborative mechanisms of rapid learning in the hippocampus and long-term integration in the neocortex [142, 143, 144, 145, 146], continual learning methods can be systematically categorized into the following three approaches.
@@
Architecture-based methods mimic the brain’s modular organization [147, 148, 149, 150] through dynamic network expansion or modular design to isolate task-specific knowledge. These methods typically encapsulate such knowledge within distinct network modules. As model sizes grow, existing approaches increasingly adopt Parameter-Efficient Fine-Tuning (PEFT) techniques (e.g., LoRA [151] and Prompt tuning [152]) to store and adapt knowledge. Formally, let $\theta_{old}$ denote the model parameters that remain frozen during training on the current task. These typically include the backbone network and the extended sub-network associated with previous tasks. The parameters specific to the current task are denoted by $\theta_{new}$, which are usually the only trainable components. At this point, Eq (1) is rewritten as follows:
$\mathop{\arg\min}_{\theta_{new}}~{}\mathbb{E}_{(x,y)\sim\mathcal{D}_{t}}\left[\ell(f_{\theta_{old}\bigcup\theta_{new}}(x),y))\right].$ (8)
@@
Since the generative model inherently possesses strong generalization capabilities, these sub-modules with relatively few parameters can effectively retain task-specific knowledge. During inference, one of the key challenges for this class of methods is to effectively select the appropriate sub-module based on the input sample.
@@
Regularization-based methods emulate neocortical synaptic stability and metaplasticity by constraining critical parameters to preserve learned representations, mirroring biological selective consolidation [153, 154, 155, 156, 157]. To achieve this goal, generative models often employ regularization in either the parameter space or the feature space, with the former formalized as follows:
@@
$\mathop{\arg\min}_{\theta}~{}\mathbb{E}_{(x,y)\sim\mathcal{D}_{t}}\left[\ell(f_{\theta}(x),y))+\lambda\cdot\Omega(\theta,\theta_{old})\right],$ (9)
@@
$\mathop{\arg\min}_{\theta}~{}\mathbb{E}_{(x,y)\sim\mathcal{D}_{t}}\left[\ell(f_{\theta}(x),y))+\lambda\cdot\mathcal{R}(\phi,\phi_{old})\right],$ (10)
where $\phi=f_{\theta}(x)$ denotes the output features or logits extracted by the model, and $\mathcal{R}(\cdot,\cdot)$ is a feature regularization term (e.g., knowledge distillation [158]) used to preserve the representations learned from previous tasks. The central challenge of such methods is to devise regularization terms that effectively retain knowledge from prior tasks with minimal impact on the acquisition of new knowledge.
Replay-based methods replicate hippocampal memory replay mechanisms [159, 160, 161, 162] by storing raw data or generating synthetic samples to revisit past experiences during new task training, effectively mitigating catastrophic forgetting. In practice, this approach typically involves maintaining an auxiliary memory buffer $\mathcal{M}$ that stores a subset of data from previous tasks, which is then used in joint training with current task data:
$\mathop{\arg\min}_{\theta}~{}\mathbb{E}_{(x,y)\sim(\mathcal{D}_{t}\bigcup\mathcal{M})}\left[\ell(f_{\theta}(x),y))\right].$ (11)
@@
3 Continual Learning in Large Language Models
In recent years, Large Language Models (LLMs) [12, 163, 164] have demonstrated outstanding natural language understanding and generation capabilities by pre-training on massive general-domain text corpora. To sustain high performance in real-world applications, LLMs require the capability to integrate new information and retain prior knowledge as data, tasks, and user preferences evolve. Unlike traditional language models, LLMs treat every task as a generative sequence prediction problem through autoregressive next-token prediction. Their massive parameterization enables emergent capabilities including few-shot learning [165] and complex multi-step reasoning [166]. Consequently, many conventional continual learning methods designed for smaller models struggle to adapt effectively to LLMs.
@@ 3 Continual Learning in Large Language Models
3.1 Problem Setup
3.1.1 Training objectives
In the instruction tuning paradigm for a LLM $f_{\theta}$, the training dataset $\mathcal{D}=\{(x_{j},y_{j})\}_{j=1}^{N}$ typically consists of input texts $x$ (e.g.,, questions, instructions, or context) paired with target output texts $y$ (e.g.,, answers, summaries, or responses), where $N$ denotes the total number of samples. Given an input $x$ and its corresponding output $y$ of length $L$, the model generates $y$ autoregressively:
$p(y\mid x)=\prod_{i=1}^{L}p_{\boldsymbol{\theta}}(y_{i}\mid x,y_{<i}),$ (12)
@@ 3.1.1 Training objectives
$\min_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta})=\mathbb{E}_{(\boldsymbol{x},\boldsymbol{y})\sim\mathcal{D}}\left[\ell(f_{\boldsymbol{\theta}}(\boldsymbol{x}),\boldsymbol{y})\right],$ (13)

### v4

@@ 2.2. Evaluation Metrics
(§ 5.3), color=cyan!100, fill=cyan!15, very thick, text=black, text width=2.5cm [CAMA (Kim et al., 2024b), M2Distill (Roy et al., 2024), color=cyan!100, fill=cyan!15, very thick, text=black, text width=7.25cm] ] [Replay-based
(§ 5.4), color=cyan!100, fill=cyan!15, very thick, text=black, text width=2.5cm [iManip (Zheng et al., 2025c), RWLA (Yang et al., 2024b), color=cyan!100, fill=cyan!15, very thick, text=black, text width=7.25cm] ] ] [Diffusion Models
(§ 6), color=darkpastelgreen!100, fill=darkpastelgreen!15, very thick, text=black, text width=2.5cm [Architecture-based
(§ 6.2), color=darkpastelgreen!100, fill=darkpastelgreen!15, very thick, text=black, text width=2.5cm [C-LoRA (Smith et al., 2023), color=darkpastelgreen!100, fill=darkpastelgreen!15, very thick, text=black, text width=7.25cm] ] [Regularization-based
(§ 6.3), color=darkpastelgreen!100, fill=darkpastelgreen!15, very thick, text=black, text width=2.5cm [MuseumMaker (Liu et al., 2024b), LFS-Diffusion (Song et al., 2024), CIDM (Dong et al., 2024), ConceptGuard (Guo and Jin, 2025),
Jha et al (Jha et al., 2025), color=darkpastelgreen!100, fill=darkpastelgreen!15, very thick, text=black, text width=7.25cm] ] [Replay-based
(§ 6.4), color=darkpastelgreen!100, fill=darkpastelgreen!15, very thick, text=black, text width=2.5cm $\text{L}^{2}\text{DM}$ (Sun et al., 2024), color=darkpastelgreen!100, fill=darkpastelgreen!15, very thick, text=black, text width=7.25cm] ] ] ]
Figure 3. Taxonomy of Continual Learning for Generative AI.
2.3. Taxonomy
@@
Inspired by the collaborative mechanisms of rapid learning in the hippocampus and long-term integration in the neocortex (Jeong et al., 2025; Moscovitch et al., 2016; Sirota et al., 2003; Kumaran et al., 2016; McClelland and Goddard, 1996), continual learning methods can be systematically categorized into the following three approaches.
@@
Architecture-based methods mimic the brain’s modular organization (Aso and Rubin, 2016; Aso et al., 2014; Frankland and Bontempi, 2005; Euston et al., 2012) through dynamic network expansion or modular design to isolate task-specific knowledge. These methods typically encapsulate such knowledge within distinct network modules. As model sizes grow, existing approaches increasingly adopt Parameter-Efficient Fine-Tuning (PEFT) techniques (e.g., LoRA (Hu et al., 2022) and Prompts (Lester et al., 2021)) to store and adapt knowledge. Formally, let $\theta_{old}$ denote the model parameters that remain frozen during training on the current task. These typically include the backbone network and the extended sub-network associated with previous tasks. The parameters specific to the current task are denoted by $\theta_{new}$, which are usually the only trainable components. At this point, Eq (1) is rewritten as follows:
(8) $\mathop{\arg\min}_{\theta_{new}}~\mathbb{E}_{(x,y)\sim\mathcal{D}_{t}}\left[\ell(f_{\theta_{old}\bigcup\theta_{new}}(x),y))\right].$
@@
Since the generative model inherently possesses strong generalization capabilities, these submodules with relatively few parameters can effectively retain task-specific knowledge. During inference, one of the key challenges for this class of methods is to effectively select the appropriate sub-module based on the input sample.
@@
Regularization-based methods emulate neocortical synaptic stability and metaplasticity by constraining critical parameters to preserve learned representations, mirroring biological selective consolidation (Hayashi-Takagi et al., 2015; Yang et al., 2009; Arevian et al., 2008; Flesch et al., 2022; Xie et al., 2022). To achieve this goal, generative models often employ regularization in either the parameter space or the feature space, with the former formalized as follows:
@@
(9) $\mathop{\arg\min}_{\theta}~\mathbb{E}_{(x,y)\sim\mathcal{D}_{t}}\left[\ell(f_{\theta}(x),y))+\lambda\cdot\Omega(\theta,\theta_{old})\right],$
@@
(10) $\mathop{\arg\min}_{\theta}~\mathbb{E}_{(x,y)\sim\mathcal{D}_{t}}\left[\ell(f_{\theta}(x),y))+\lambda\cdot\mathcal{R}(\phi,\phi_{old})\right],$
where $\phi=f_{\theta}(x)$ denotes the output features or logits extracted by the model, and $\mathcal{R}(\cdot,\cdot)$ is a feature regularization term (e.g., knowledge distillation (Hinton et al., 2015)) used to preserve the representations learned from previous tasks. The central challenge of such methods is to devise regularization terms that effectively retain knowledge from prior tasks with minimal impact on the acquisition of new knowledge.
Replay-based methods replicate hippocampal memory replay mechanisms (Davidson et al., 2009; Gupta et al., 2010; Carr et al., 2011; Ólafsdóttir et al., 2018) by storing raw data or generating synthetic samples to revisit past experiences during new task training, effectively mitigating catastrophic forgetting. In practice, this approach typically involves maintaining an auxiliary memory buffer $\mathcal{M}$ that stores a subset of data from previous tasks, which is then used in joint training with current task data:
(11) $\mathop{\arg\min}_{\theta}~\mathbb{E}_{(x,y)\sim(\mathcal{D}_{t}\bigcup\mathcal{M})}\left[\ell(f_{\theta}(x),y))\right].$
@@
3. Continual Learning for Large Language Models
In recent years, Large Language Models (LLMs) (Achiam et al., 2023; Grattafiori et al., 2024; Yang et al., 2024c) have demonstrated outstanding natural language understanding and generation capabilities by pre-training on massive general-domain text corpora. To sustain high performance in real-world applications, LLMs require the capability to integrate new information and retain prior knowledge as data, tasks, and user preferences evolve. Unlike traditional language models, LLMs treat every task as a generative sequence prediction problem through autoregressive next-token prediction. Their massive parameterization enables emergent capabilities including few-shot learning (Brown et al., 2020) and complex multi-step reasoning (Wei et al., 2022). Consequently, many conventional continual learning methods designed for smaller models struggle to adapt effectively to LLMs.
@@ 3. Continual Learning for Large Language Models
3.1. Problem Setup
Training objectives. In the instruction tuning paradigm for a LLM $f_{\theta}$, the training dataset $\mathcal{D}=\{(x_{j},y_{j})\}_{j=1}^{N}$ typically consists of input texts $x$ (e.g., questions) paired with target output texts $y$ (e.g., answers), where $N$ denotes the total number of samples. Given an input $x$ and its corresponding output $y$ of length $L$, the model generates $y$ autoregressively:
(12) $p(y\mid x)=\prod_{i=1}^{L}p_{\boldsymbol{\theta}}(y_{i}\mid x,y_{<i}),$
@@ 3.1. Problem Setup
(13) $\min_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta})=\mathbb{E}_{(\boldsymbol{x},\boldsymbol{y})\sim\mathcal{D}}\left[\ell(f_{\boldsymbol{\theta}}(\boldsymbol{x}),\boldsymbol{y})\right],$